You have purchased a bond 30 months ago with par=$1000, semi-annual coupons of 30 (=PMT) and YTM of 4%.

If you sell the bond today, what is the value of coupons you have received while holding the bond (i.e.FVcoupons) today, given a coupon reinvestment rate of 7% over the past 30 months?

Please explain solution and please be sure of the work.
Answer #1

Semi annual coupon =30
Semi Reinvestment rate =7%/2 =3.5%
Number of periods =30/6 =5
FV of Coupons =Coupon*((1+r)^n-1)/r=30*((1+3.5%)^5-1)/3.5%) =160.87